A doctor's age 8 years ago is two-thirds his age 13 years hance. How old is he now?
Answer by anamika1310(3)

let the doctor's age be x.
so,
x-8 = 2/3(x+13)
3x-24=2x+26
x=50
So, his present age is 50.
hope this is it.